y' = csc x - y cot x


1
Expert's answer
2021-10-17T17:46:07-0400

Let us solve the differential equation "y' = \csc x - y \cot x," which is equivalent to "y' + y \frac{\cos x}{\sin x}= \frac{1}{\sin x}."

Let us multiply both part of the equation by "\sin x." Then we get the equation "y'\sin x+y\cos x=1," which is equivalent to "(y\sin x)'=1." It follows that "y\sin x= x+C," and hence the general solution is of the form

"y=\frac{1}{\sin x}(x+C)=\csc x(x+C)."